# EP241021a: A catastrophic collapse/merger of compact star binary leading to the formation of a remnant millisecond magnetar?

### v1

@@ 3.2 The external shock of a jet
The existence of the bright radio emission and the underlying power-law decaying component of the radio and X-ray emission strongly indicate that a powerful jet could be driven by the WD collapses or DNS/NS-WD mergers, which gives rise to multi-wavelength non-thermal emission through an external shock by interacting with the surrounding medium. On the one hand, the relativistic degree of this jet could not be as high as those of normal GRBs, at least on the viewing direction, because no GRB emission was detected accompanying the FXRT emission. On the other hand, the jet could in principle have a complicated angular structure (e.g., as conceived in Gianfagna et al. 2025), which would be beneficial for explaining the detailed features appearing in the observed radio light curves. Nevertheless, in this Letter, we take into account a top-hat jet simply, which would be found to be good enough to describe the general evolution behavior of the underlying non-thermal emission in all of the radio, optical, and X-ray bands.
The dynamical evolution of a relativistic jet of a bulk Lorentz factor $\Gamma_{\rm j}$ can be described by (Blandford & McKee, 1976; Huang et al., 1999)
@@
where $E_{\mathrm{j}}$ is the isotropically equivalent kinetic energy of the jet, $n_{0}$ is the particle number density of the interstellar medium, and $m_{\mathrm{p}}$ is the proton mass. The electrons, which are accelerated by the external shock driven by the jet, would be distributed on its comoving random Lorentz factor as a power law $\gamma^{-p}$ with a minimum value of $\gamma_{\mathrm{m}}=\epsilon_{\mathrm{e}}(\Gamma_{\mathrm{j}}-1)\left(p-2\right)m_{\mathrm{p}}/[\left(p-1\right)m_{\mathrm{e}}]$, where $\epsilon_{\mathrm{e}}$ is the equipartition factor for the electrons. Considering the modification of the electron distribution due to their synchrotron cooling, a cooling Lorentz factor is usually defined as $\gamma_{\mathrm{c}}=(1+z)6\pi m_{\mathrm{e}}c/\sigma_{T}B_{\mathrm{j}}^{\prime 2}\Gamma_{\mathrm{j}}t$, where $B^{\prime}_{\mathrm{j}}=\sqrt{32\pi\epsilon_{B}\Gamma_{\mathrm{j}}^{2}n_{0}m_{\mathrm{p}}c^{2}}$ is the strength of the stochastic magnetic field with $\epsilon_{B}$ being its equipartition factor.
@@
$L_{\mathrm{j},\nu}=$ $L_{\mathrm{j,max}}\frac{(1-e^{-\tau_{\mathrm{ssa}},\nu})}{\tau_{\mathrm{ssa},\nu}}$ (8)
×$\times\begin{cases}\left(\frac{\nu}{\nu_{\mathrm{l}}}\right)^{1/3},&\nu<\nu_{\mathrm{l}};\\[10.0pt] \left(\frac{\nu}{\nu_{\mathrm{l}}}\right)^{-(q-1)/2},&\nu_{\mathrm{l}}<\nu<\nu_{\mathrm{h}};\\[10.0pt] \left(\frac{\nu_{\mathrm{h}}}{\nu_{\mathrm{l}}}\right)^{-(q-1)/2}\left(\frac{\nu}{\nu_{\mathrm{h}}}\right)^{-p/2},&\nu_{\mathrm{h}}<\nu,\end{cases}$
where $\nu_{\mathrm{l}}=\mathrm{min}\left[\nu_{\mathrm{m}},\nu_{\mathrm{c}}\right]$ and $\nu_{\mathrm{h}}=\mathrm{max}\left[\nu_{\mathrm{m}},\nu_{\mathrm{c}}\right]$, $q=2$ for $\nu_{\mathrm{m}}>\nu_{\mathrm{c}}$ and $q=p$ for $\nu_{\mathrm{m}}<\nu_{\mathrm{c}}$. The spectral peak luminosity reads $L_{\mathrm{j,max}}=(1+z)\Gamma_{\rm j}N_{\mathrm{e}}m_{\mathrm{e}}c^{2}\sigma_{T}B^{\prime}_{\mathrm{j}}/3q_{\mathrm{e}}$, where $m_{\mathrm{e}}$, $q_{\mathrm{e}}$ and $\sigma_{\mathrm{T}}$ are the electron mass, charge, and the Thomson cross-section, respectively. The total number of electrons is given by $N_{\mathrm{e}}=4\pi R_{\mathrm{j}}^{3}n_{0}/{3}$ with $R_{\mathrm{j}}=[17E_{\mathrm{j}}t/4\pi(1+z)m_{\mathrm{p}}n_{0}c]^{1/4}$. Finally, the synchrotron self-absorption depth can be expressed as
$\tau_{\mathrm{ssa},\nu}=$ $\frac{q_{\mathrm{e}}N_{\mathrm{e}}}{4\pi B^{\prime}_{\mathrm{j}}R_{{\mathrm{j}}}^{2}\gamma_{\mathrm{l}}^{5}}$ (9)
×$\times\begin{cases}\left(\frac{\nu}{\nu_{\mathrm{l}}}\right)^{-5/3},&\nu<\nu_{\mathrm{l}};\\[10.0pt] \left(\frac{\nu}{\nu_{\mathrm{l}}}\right)^{-(q+4)/2},&\nu_{\mathrm{l}}<\nu<\nu_{\mathrm{h}};\\[10.0pt] \left(\frac{\nu_{\mathrm{h}}}{\nu_{\mathrm{l}}}\right)^{-(q+4)/2}\left(\frac{\nu}{\nu_{\mathrm{h}}}\right)^{-(q+5)/2},&\nu_{\mathrm{h}}<\nu.\end{cases}$
where $\gamma_{\mathrm{{l}}}=\mathrm{min}[\gamma_{\mathrm{m}},\gamma_{\mathrm{c}}]$.
@@ 4 Light curve fittings
In order to confront with the observational data, the flux density of the optical emission at observed frequency $\nu$ can be given by $F_{\nu}=L_{\nu}/4\pi D_{L}^{2}$, where $D_{L}$ is the luminosity distance. The apparent magnitude is given by $M_{\nu}=-2.5~{}\mathrm{log_{10}}\left(F_{\nu}\right)-48.6$. By comparison, the generation of the X-ray emission could be most complicated. As illustrated in Figure 4, besides the contribution from the jet, the emission leaking from the MWN and the magnetic reconnection zone could gradually become the dominant component of the X-ray emission, which is further direction-dependent.
@@
On the one hand, the MWN emission itself could be anisotropic, so we introduce a free parameter $\zeta$ to represent the fraction of the spin-down energy released into the jet direction. If $\zeta\rightarrow 1$, then the energy is highly beamed in the jet direction. In this case, an ultra-relativistic jet could be easily launched. Otherwise, the energy is anti-beamed for $\zeta\ll\Omega_{\rm j}/4\pi$, which could be unfavorable to the formation of an ultra-relativistic jet, where $\Omega_{\rm j}$ is the solid opening angle of the jets. On the other hand, Moreover, the leakage of the MWN emission is further dependent on the anisotropic structure of the jet and ejecta. In the jet direction, the MWN emission is likely to be detectable from the very moment of the beginning, because the jet material is always transparent. On the contrary, in the other directions, the MWN emission can be detected only after the explosion ejecta becomes optical thin for X-rays. Following the above considerations, the total X-ray luminosity can be expressed as
@@
where the X-ray luminosity of the MWN is simply assumed to trace the spin-down luminosity of the magnetar without a detailed description of the MWN emission mechanisms. $f_{\rm j}=\Omega_{\rm j}/4\pi$ is introduced because the MWN at the jet direction is probably relativistic. Here we take $f_{\rm j}\approx 0.15$ as a reference value, roughly corresponding to a half-opening angle around $30^{\circ}$ of the jet.
@@
$E_{\rm j}/10^{51}\mathrm{erg}$ Log-flat [0.01,100] $3.01_{-0.04}^{+0.04}$
$n_{0}/\mathrm{cm^{-3}}$ Log-flat [0.001,10] $0.16_{-0.03}^{+0.03}$
$p$ Flat [2,3] $2.35_{-0.01}^{+0.01}$
$\epsilon_{\mathrm{e}}$ Flat [0.01,1] $0.12_{-0.00}^{+0.00}$
$\epsilon_{B}$ Log-flat [0.0001,0.1] $0.04_{-0.01}^{+0.01}$
@@
$\kappa/\mathrm{cm^{2}g^{-1}}$ Flat [0.1,50] $12.43_{-4.55}^{+4.51}$
$\kappa_{\mathrm{x}}/\mathrm{cm^{2}g^{-1}}$ Flat [0.1,10] $1.79_{-0.67}^{+0.68}$
$T_{\mathrm{floor}}/10^{3}\mathrm{K}$ Flat [1,10] $8.18_{-0.19}^{+0.20}$
@@
$P_{\mathrm{m}}/\mathrm{ms}$ Flat [0.5,10] $5.41_{-0.14}^{+0.13}$
$B_{\mathrm{m}}/10^{14}\mathrm{G}$ Flat [0.1,10] $3.45_{-0.11}^{+0.12}$
$\zeta$ Log-Flat [0.001,1] $0.04_{-0.00}^{+0.00}$
@@
We perform the light curve fittings for EP241021a with the Markov Chain Monte Carlo method of 50000 iterations, by using the emcee package (Foreman-Mackey et al., 2013). There are five free parameters for the jet emission and eight free parameters relevant to the emission powered by the magnetar. The log likelihood function is $\mathrm{ln}~{}\mathcal{L}=-0.5(\Sigma_{i=1}^{n}(O_{i}-M_{i})^{2}/\sigma_{i}^{2})$, where $O_{i}$, $M_{i}$ and $\sigma_{i}$ are the $i-\mathrm{th}$ observed flux densities or apparent magnitudes, the model outputs, and the observational uncertainties, respectively. The fitting results are shown by the lines in Figures 1, 2 and 3 for the X-ray, optical, and radio light curves, respectively. The priors and posteriors with $16\%$ and $84\%$ quantiles are summarized in Table 1. The corner plot of the posterior probability distributions is presented in Figure 5. Furthermore, in Figures 6 and 7, we further show the general consistency between the model and the optical and radio data in the form of a group of evolving spectra. It should be mentioned that, around the peak time $T_{0}+10$ days, a relatively prominent excess could exist on the red side of the spectra, which can also be found in the light curve fittings.
Generally, our model provides a reasonable explanation for the multi-wavelength observations of EP241021a, including all of the X-ray, optical, and radio emission. In particular, the derived parameters for the ejecta agree well with our analyses proposed in Section 3. The mass, speed, and opacity of the ejecta all direct it towards the origin of a catastrophic compact star collapse/merger. The small value of $\zeta$, which is probably smaller than $f_{\rm j}$, further indicates that the energy release from the magnetar could be somewhat anti-beamed in the polar direction. The rough equivalence between the isotropic kinetic energies of the jet and the explosion ejecta further strengthens the fact that the magnetar released energy roughly isotropically. This may be a reason why no GRB emission can be detected from the EP241021a event.
@@ 5 Conclusion and Discussion
The newly discovered EP241021a was featured by a multi-wavelength rebrightening around ten days after the WXT trigger. The appearance of the X-ray emission indicates that this rebrightening cannot be contributed by an interaction-powered SN [^1] If only the optical emission was detected, then in principle one cannot rule out that the significant optical bump is contributed by the cooling of SN shock-heated material, because the non-detection of SN signal could just be due to the large distance of the source. Although the bolometric luminosity and peak timescale of the optical bump are analogous to FBOTs, its relatively red color still indicates it as a new transient phenomenon. The primary common feature between EP241021a and FBOTs as well as GRBs could be the existence of a powerful magnetar engine, which may provide us a clue to approach a unified understanding of various extreme stellar explosion phenomena (Yu et al., 2017; Liu et al., 2022). As a specific example, EP241021a could hint that the energy release from the rapidly rotating magnetar could not always be collimated in the polar direction on a degree as high as in the GRB case. The anisotropic degree of the energy release of the central engine could be a crucial factor determining the relativistic degree of the jet. Meanwhile, the anisotropic structure of the explosion ejecta would also play an important role in shaping the details of the spectra and light curves of the optical transients, e.g., leading to an obvious excess in the $J$ band.
@@
The reason why EP241021a did not produce GRB radiation but only generated an FXRT might be that the relativistic degree of the jet was not high enough, or it could be that the internal variability in the jet was not significant enough to cause sufficient internal dissipation. It is actually difficult to determine whether the specific mechanism causing FXRT emission is the same as that of GRB, with only differences in energy bands. If so, the duration of EP241021a may tend to support the origins of NS-WD merger and WD collapse, as a DNS merger hardly maintains such a long accretion duration. Furthermore, on the one hand, in contrast to the WD collapse, the NS-WD merger could have an advantage in forming a millisecond magnetar and driving a moderate relativistic jet. On the other hand, if the ambiguous second bump appearing at around $T_{0}+50$ days is real, then the accretion-induced collapse (AIC) of a WD would become most favored, because the millisecond magnetar could evaporate the residual companion star quasi-periodically due to their orbital motion. As a result, this evaporated material would provide delayed energy injection into the AIC ejecta, leading to post-peak bumps in the light curve of the ejecta emission (Zhu et al., 2024). In any case, the advantage of the DNS is that it is most likely to eject sufficient neutron-rich material to explain the high opacity required by the observation. If the prompt X-ray emission is not contributed by the jet, then this possibility cannot be ruled out. If so, EP241021a might be regarded as the first luminous mergernova candidate that was suggested by (Yu et al., 2013).
@@
As proposed by Zhang (2013), FXRT emission could in principle directly come from the relativistic wind of a magnetar. This has previously been used to explain some observed X-ray orphan transients such as CDF-S XT2 (Xue et al., 2019) and some prompt X-ray emission during GRBs (e.g., GRB 230307A; Sun et al., 2025). If the prompt X-ray emission of EP241021a also has such an origin, then its 100s-duration may correspond to an early spin-down timescale of the magnetar. Here, one may note that the fitting of the optical bump has determined $t_{\rm sd}=5.8$ days, which seems to contradict the above hypothesis. However, the critical constraint on the magnetar parameters from the optical fitting can actually be interpreted as follows: the magnetar should have a spin-down luminosity of $L_{\rm sd,i}=1.4\times 10^{45}\rm erg~{}s^{-1}$ around the time $t_{\rm sd}=5.8$ days. For timescales much shorter than 5.8 days, the optical data cannot provide a serious constraint on whether the spin-down luminosity is the same as at 5.8 days or significantly higher. In other words, it can not be ruled out that the magnetar may have experienced continuous deceleration within the first several days rather than maintaining a constant spin as assumed. In the GRB 230307A case, it was found that the prompt X-ray emission and the following magnetar-powered kilonova-like emission can be easily connected by a unified spin-down behavior (Wang et al., 2024). However, for EP241021a, since the luminosity of the optical bump is extremely high, it seems difficult to connect the prompt X-ray emission and the optical bump by a simple magnetic braking process. This may also indicate EP241021a could have an origin different from GRB 230307A which is likely to be produced by an NS-WD merger.
@@
In any case, if more elaborated modeling is sought in the future, a more complicated braking mechanism could be taken into account. Moreover, the dynamical evolution of the magnetar wind and its specific emission processes need to be described more carefully (e.g., Yu et al., 2019), so that a more realistic X-ray light curve can be obtained. For example, the abrupt decay of the X-ray emission at $T_{0}+80$ days could be explained by the shift of the emission from X-ray to lower-energy band. Finally, an elaborate description of the ejecta structure and a frequency-dependent opacity would also help to understand the detailed light curve and spectral characteristics.

### v2

@@ 3.2 The external shock of a jet
The existence of the bright radio emission and the underlying power-law decaying component of the radio and X-ray emission strongly indicate that a powerful jet could be driven by the WD collapses or DNS/NS-WD mergers, which gives rise to multi-wavelength non-thermal emission through an external shock by interacting with the surrounding medium. On the one hand, the relativistic degree of this jet could not be as high as those of normal GRBs, at least on the viewing direction, because no GRB emission was detected accompanying the FXRT emission. On the other hand, the jet could in principle have a complicated angular structure (e.g., as conceived in Gianfagna et al. 2025), which would be beneficial for explaining the detailed features appearing in the observed radio light curves. Nevertheless, in this paper, we take into account a top-hat jet simply, which would be found to be good enough to describe the general evolution behavior of the underlying non-thermal emission in all of the radio, optical, and X-ray bands.
The interaction of a relativistic jet with the surrounding medium would initially drive a reverse shock propagating into the jet and a forward shock propagating into the medium. After the reverse shock crosses the whole jet during the first few hundreds of seconds, the forward shock (i.e., a blast wave) would enter into the Blandford-McKee self-similar evolution phase, during which the bulk Lorentz factor of the shock can be written as (Blandford & McKee, 1976; Huang et al., 1999):
@@
where $E_{\mathrm{j}}$ is the isotropically equivalent kinetic energy of the jet, $n_{0}$ is the particle number density of the interstellar medium, and $m_{\mathrm{p}}$ is the proton mass. The electrons, which are accelerated by the forward shock, would be distributed on its comoving random Lorentz factor as a power law $\gamma^{-p}$ with a minimum value of $\gamma_{\mathrm{m}}=\epsilon_{\mathrm{e}}(\Gamma_{\mathrm{j}}-1)\left(p-2\right)m_{\mathrm{p}}/[\left(p-1\right)m_{\mathrm{e}}]$, where $\epsilon_{\mathrm{e}}$ is the equipartition factor for the electrons. Considering the modification of the electron distribution due to their synchrotron cooling, a cooling Lorentz factor is usually defined as $\gamma_{\mathrm{c}}=(1+z)6\pi m_{\mathrm{e}}c/\sigma_{T}B_{\mathrm{j}}^{\prime 2}\Gamma_{\mathrm{j}}t$, where $B^{\prime}_{\mathrm{j}}=\sqrt{32\pi\epsilon_{B}\Gamma_{\mathrm{j}}^{2}n_{0}m_{\mathrm{p}}c^{2}}$ is the strength of the stochastic magnetic field with $\epsilon_{B}$ being its equipartition factor.
@@
$L_{\mathrm{j},\nu}=$ $L_{\mathrm{j,\nu,max}}\frac{(1-e^{-\tau_{\mathrm{ssa}},\nu})}{\tau_{\mathrm{ssa},\nu}}$ (8)
[…]
where $\nu_{\mathrm{l}}=\mathrm{min}\left[\nu_{\mathrm{m}},\nu_{\mathrm{c}}\right]$ and $\nu_{\mathrm{h}}=\mathrm{max}\left[\nu_{\mathrm{m}},\nu_{\mathrm{c}}\right]$, $q=2$ for $\nu_{\mathrm{m}}>\nu_{\mathrm{c}}$ and $q=p$ for $\nu_{\mathrm{m}}<\nu_{\mathrm{c}}$. The spectral peak luminosity reads $L_{\mathrm{j,\nu,max}}=(1+z)\Gamma_{\rm j}N_{\mathrm{e}}m_{\mathrm{e}}c^{2}\sigma_{T}B^{\prime}_{\mathrm{j}}/3q_{\mathrm{e}}$,
where $m_{\mathrm{e}}$, $q_{\mathrm{e}}$ and $\sigma_{\mathrm{T}}$ are the electron mass, charge, and the Thomson cross-section, respectively. The total number of electrons is given by $N_{\mathrm{e}}=4\pi R_{\mathrm{j}}^{3}n_{0}/{3}$ with $R_{\mathrm{j}}=[17E_{\mathrm{j}}t/4\pi(1+z)m_{\mathrm{p}}n_{0}c]^{1/4}$. Finally, the synchrotron self-absorption depth can be expressed as
$\tau_{\mathrm{ssa},\nu}=$ $C\frac{q_{\mathrm{e}}N_{\mathrm{e}}}{4\pi B^{\prime}_{\mathrm{j}}R_{{\mathrm{j}}}^{2}\gamma_{\mathrm{l}}^{5}}$ (9)
[…]
where $\gamma_{\mathrm{{l}}}=\mathrm{min}[\gamma_{\mathrm{m}},\gamma_{\mathrm{c}}]$ and $C\approx 5$ is a factor slightly depending on $p$ (Panaitescu & Mészáros, 2000).
@@ 4 Light curve fittings
In order to confront with the observational data, the flux density of the optical emission at observed frequency $\nu$ can be given by $F_{\nu}=L_{\nu}/4\pi D_{L}^{2}$, where $D_{L}$ is the luminosity distance. The apparent magnitude is given by $M_{\nu}=-2.5~\mathrm{log_{10}}\left(F_{\nu}\right)-48.6$ in the AB system where $F_{\nu}$ is in units of $\mathrm{erg~s^{-1}cm^{-2}Hz^{-1}}$. By comparison, the generation of the X-ray emission could be most complicated. As illustrated in Figure 4, besides the contribution from the jet, the emission leaking from the MWN and the magnetic reconnection zone could gradually become the dominant component of the X-ray emission, which is further direction-dependent. In our calculations, the observer is assumed on the axis of the jet.
@@
On the one hand, the MWN emission itself could be anisotropic, so we introduce a free parameter $\zeta$ to represent the fraction of the spin-down energy released into the jet direction. If $\zeta\rightarrow 1$, then the energy is highly beamed in the jet direction. In this case, an ultra-relativistic jet could be easily launched. Otherwise, the energy is anti-beamed for $\zeta\ll\Omega_{\rm j}/4\pi$, which could be unfavorable to the formation of an ultra-relativistic jet, where $\Omega_{\rm j}$ is the solid opening angle of the jets. On the other hand, Moreover, the leakage of the MWN emission is further dependent on the anisotropic structure of the jet and ejecta. In the jet direction, the MWN emission is likely to be detectable from the very moment of the beginning, because the jet material is always transparent. On the contrary, in the other directions, the MWN emission can be detected only after the explosion ejecta becomes optical thin for X-rays. Following the above considerations, the total luminosity in X-ray band can be expressed as
@@
where the X-ray luminosity of the MWN is simply assumed to trace the spin-down luminosity of the magnetar without a detailed description of the MWN emission mechanisms. $f_{\rm j}=\Omega_{\rm j}/4\pi$ is introduced because the MWN at the jet direction is probably relativistic and, consequently, its emission could be collimated within such a solid angle rather than isotropic. Here we take $f_{\rm j}\approx 0.15$ as a reference value, roughly corresponding to a half-opening angle around $30^{\circ}$ of the jet.
@@
$E_{\rm j}/10^{51}\mathrm{erg}$ Log-flat [0.01,100] $2.37_{-0.17}^{+0.18}$
$n_{0}/\mathrm{cm^{-3}}$ Log-flat [0.001,10] $0.28_{-0.07}^{+0.10}$
$p$ Flat [2,3] $2.34_{-0.01}^{+0.01}$
$\epsilon_{\mathrm{e}}$ Flat [0.01,1] $0.15_{-0.01}^{+0.01}$
$\epsilon_{B}$ Log-flat [0.0001,0.1] $0.03_{-0.01}^{+0.01}$
@@
$\kappa/\mathrm{cm^{2}g^{-1}}$ Flat [0.1,50] $12.58_{-4.52}^{+3.93}$
$\kappa_{\mathrm{x}}/\mathrm{cm^{2}g^{-1}}$ Flat [0.1,10] $1.83_{-0.65}^{+0.63}$
$T_{\mathrm{floor}}/10^{3}\mathrm{K}$ Flat [1,10] $8.19_{-0.19}^{+0.20}$
@@
$L_{\rm sd,i}/\mathrm{10^{47}~erg~s^{-1}}$ Log-Flat [0.001,1000] $0.01_{-0.002}^{+0.002}$
$t_{\mathrm{sd}}/10^{5}\rm s$ Log-Flat [0.001,1000] $5.19_{-0.51}^{+0.52}$
$\zeta$ Log-Flat [0.001,1] $0.04_{-0.005}^{+0.005}$
@@
We perform the light curve fittings for EP241021a with the Markov Chain Monte Carlo method of 50000 iterations, by using the emcee package (Foreman-Mackey et al., 2013). There are five free parameters for the jet emission and eight free parameters relevant to the emission powered by the magnetar. The log likelihood function is $\mathrm{ln}~\mathcal{L}=-0.5(\Sigma_{i=1}^{n}(O_{i}-M_{i})^{2}/\sigma_{i}^{2})$, where $O_{i}$, $M_{i}$ and $\sigma_{i}$ are the $i-\mathrm{th}$ observed flux densities or apparent magnitudes, the model outputs, and the observational uncertainties, respectively. The fitting results are shown by the lines in Figures 1, 2 and 3 for the X-ray, optical, and radio light curves, respectively. The priors and posteriors with $16\%$ and $84\%$ quantiles are summarized in Table 1. The corner plot of the posterior probability distributions is presented in Figure 5. Furthermore, in Figures 6, we further show the general consistency between the model and the optical data in the form of a group of evolving spectra. This indicates that the black body assumption can indeed work for describing the optical bump emission, although its uniqueness for individual spectra should not be overstated. In addition, one may note that, around the peak time $T_{0}+10$ days, a relatively prominent excess could exist on the red side of the spectra, which can also be found in the light curve fittings. This reveals the complexity of the spectral characteristics of this source. Finally, comparisons between the observational radio spectra and the synchrotron spectra of the jet external shock emission are presented in Figure 7.
Generally, our model provides a reasonable explanation for the multi-wavelength observations of EP241021a, including all of the X-ray, optical, and radio emission. It should be noted that we focus on the global evolutionary trends of the multi-wavelength emission, rather than the shorter-timescale fluctuations in light curves. These fluctuations, potentially induced by the inhomogeneous structure of ejecta and internal local activities, are beyond the scope of our simplified model. In any case, the derived parameters for the ejecta agree well with our analyses proposed in Section 3. The mass, speed, and opacity of the ejecta all direct it towards the origin of a catastrophic compact star collapse/merger. The small value of $\zeta$, which is probably smaller than $f_{\rm j}$, further indicates that the energy release from the magnetar could be somewhat anti-beamed in the polar direction but does not deviate from isotropic very much. Even more, the rough equivalence between the isotropic kinetic energies of the jet and the explosion ejecta further strengthens the fact that the magnetar released energy roughly isotropically. This may be a reason why no GRB emission can be detected from the EP241021a event.
@@ 5 Conclusion and Discussion
The newly discovered EP241021a was featured by a multi-wavelength rebrightening around ten days after the WXT trigger. The appearance of the X-ray emission indicates that this rebrightening cannot be contributed by an interaction-powered SN [^2] If only the optical emission was detected, then in principle one cannot rule out that the significant optical bump is contributed by the cooling of SN shock-heated material, because the non-detection of SN signal could just be due to the large distance of the source.. Alternatively, the bolometric luminosity and peak timescale of the optical bump make it analogous to FBOTs, if its spectra can indeed be described by a black body, which although still needs to be confirmed. The primary common feature between EP241021a and FBOTs as well as GRBs could be the existence of a powerful magnetar engine, which may provide us a clue to approach a unified understanding of various extreme stellar explosion phenomena (Yu et al., 2017; Liu et al., 2022). As a specific example, EP241021a could hint that the energy release from the rapidly rotating magnetar could not always be collimated in the polar direction on a degree as high as in the GRB case. The anisotropic degree of the energy release of the central engine could be a crucial factor determining the relativistic degree of the jet. Meanwhile, the anisotropic structure of the explosion ejecta would also play an important role in shaping the details of the spectra and light curves of the optical transients, e.g., leading to an obvious excess in the $J$ band.
@@
The reason why EP241021a did not produce GRB radiation but only generated an FXRT might be that the relativistic degree of the jet was not high enough, or it could be that the internal variability in the jet was not significant enough to cause sufficient internal dissipation. It is actually difficult to determine whether the specific mechanism causing FXRT emission is the same as that of GRB, with only differences in energy bands. If so, the duration of EP241021a may tend to support the origins of NS-WD merger and WD collapse, as a DNS merger hardly maintains such a long accretion duration. Furthermore, on the one hand, in contrast to the WD collapse, the NS-WD merger could have an advantage in forming a millisecond magnetar and driving a moderate relativistic jet. On the other hand, if the ambiguous second bump appearing at around $T_{0}+50$ days is real, as argued by (Shu et al., 2025), then the accretion-induced collapse (AIC) of a WD would become most favored. In this scenario, as proposed by Zhu et al. (2024), the millisecond magnetar could evaporate the residual companion star quasi-periodically due to their orbital motion. As a result, this evaporated material would provide delayed energy injection into the AIC ejecta, leading to post-peak bumps in the light curve of the ejecta emission. In any case, the advantage of the DNS is that it is most likely to eject sufficient neutron-rich material to explain the high opacity required by the observation. If the prompt X-ray emission is not contributed by the jet, then this possibility cannot be ruled out. If so, EP241021a might be regarded as the first luminous mergernova candidate that was suggested by (Yu et al., 2013).
@@
As proposed by Zhang (2013), FXRT emission could in principle directly come from the relativistic wind of a magnetar. This has previously been used to explain some observed X-ray orphan transients such as CDF-S XT2 (Xue et al., 2019) and some prompt X-ray emission during GRBs (e.g., GRB 230307A; Sun et al., 2025). If the prompt X-ray emission of EP241021a also has such an origin, then its 100s-duration may correspond to an early spin-down timescale of the magnetar. Here, one may note that the fitting of the optical bump has determined $t_{\rm sd}=6$ days, which seems to contradict the above hypothesis. However, the critical constraint on the magnetar parameters from the optical fitting can actually be interpreted as follows: the magnetar should have a spin-down luminosity of $L_{\rm sd,i}=10^{45}\rm erg~s^{-1}$ around the time $t_{\rm sd}=6$ days. For timescales much shorter than 6 days, the optical data cannot provide a strong constraint on whether the spin-down luminosity is the same as at 6 days or significantly higher. In other words, it can not be ruled out that the magnetar may have experienced continuous deceleration within the first several days rather than maintaining a constant spin as assumed. In the GRB 230307A case, it was found that the prompt X-ray emission and the following magnetar-powered kilonova-like emission can be easily connected by a unified spin-down behavior (Wang et al., 2024). However, for EP241021a, since the luminosity of the optical bump is extremely high, it seems difficult to connect the prompt X-ray emission and the optical bump by a simple magnetic braking process. This may also indicate that EP241021a could have an origin different from GRB 230307A which is likely to be produced by an NS-WD merger. In more refined considerations, we need to examine whether the early energy injection might exceed the initial kinetic energy of the ejecta. If so, we must also account for the dynamical evolution of the ejecta in the model.
@@
In any case, if more elaborate modeling is sought in the future, a more complicated braking mechanism could be taken into account. The dynamical evolution of the magnetar wind and its specific emission processes also need to be described carefully (e.g., Yu et al., 2019; Wu et al., 2021), so that a refined X-ray light curve can be obtained. In this case, the abrupt decay of the X-ray emission at $T_{0}+80$ days, which is indicated by the XMM-Newton upper limits, could be explained by a spectral shift of the magnetar wind emission from the X-ray to lower-energy bands. Such a softening of the wind emission could be a natural result of the decline of the spin-down luminosity and the expansion of the wind nebula. Finally, a detailed description of the ejecta structure and frequency-dependent opacity would also help to understand the detailed light-curve and spectral characteristics.